OpenStudy (wesdg1978):

A point P in the first quadrant lies on the graph of the function f(x)=x^2. Express the slope of the line joining P to the origin as a function of x.

OpenStudy (anonymous):

the point has the coordinates as P = (x,x^2) the origin has the coordinates as O = (0,0) now the slope of the line from point P to origin = (x^2-0)/(x-0) = x^2/x = x thus slope = m = x

OpenStudy (wesdg1978):

That was my first guess but I wasn't sure if it was supposed to be set up as x^2/x or not. Thanks for your help.

OpenStudy (anonymous):

you are more than welcome